OpenStudy (anonymous):

Simplify without using a calculator. 1.) log3(log3(log3 27) 2.) e^ln 2-ln 6 3.) e^b ln a 4.) ln(3+x)-ln(x^2-9)

OpenStudy (solomonzelman):

My closest interpretation. \(\LARGE\color{blue}{ \bf log[~~~log_3(log_327)~~~] }\) \(\LARGE\color{blue}{ \bf log[~~~log_3(log_33^3)~~~] }\) \(\LARGE\color{blue}{ \bf log[~~~log_3(3log_33)~~~] }\) \(\LARGE\color{blue}{ \bf log[~~~log_3(3 \times 1)~~~] }\) \(\LARGE\color{blue}{ \bf log[~~~log_3(3 )~~~] }\) \(\LARGE\color{blue}{ \bf log[~~~1~~~] }\) \(\LARGE\color{blue}{ \bf log_{~any~~possible~~base~}1=0 }\)

OpenStudy (solomonzelman):

\(\LARGE\color{red}{ \bf Log_vv=1 }\) \(\large\color{red}{ \bf (if~~v≠0,1,complex~num.~~,~~imaginary~) }\) \(\LARGE\color{red}{ \bf Log_cc^b=b \times Log_cc=b \times 1=b }\) \(\LARGE\color{red}{ \bf Log_v1=0 }\) \(\large\color{red}{ \bf (for~~any~~value~~of~~v.) }\)

OpenStudy (solomonzelman):

questions 2 and 3 are unclear to me. I don't know how to intepret them, and 4.... \(\LARGE\color{purple }{ \bf ln(3+x)-ln(x^2-9) }\) \(\LARGE\color{purple }{ \bf ln(x+3)-ln(x^2-9) }\) \(\LARGE\color{purple }{ \bf ln \frac{(x+3)}{(x^2-9) } }\) \(\LARGE\color{purple }{ \bf ln \frac{(x+3)}{(x-3)(x+3) } }\) \(\LARGE\color{purple }{ \bf ln \frac{\cancel{ (x+3)} }{ \cancel{ (x+3)}(x-3) } }\) \(\LARGE\color{purple }{ \bf ln \frac{1}{(x-3) } }\)